OpenStudy (anonymous):

Solve the indefinite integral: e^(2x) / x^3

OpenStudy (anonymous):

where is the dx??

OpenStudy (anonymous):

on way might be to write it as e^(2x) x^(-3) and use integration by parts

OpenStudy (anonymous):

But when using integration by parts for that, wouldn't it just go on forever, nesting integration by parts?

OpenStudy (anonymous):

yup, and make u = x^-3 dv = e^2x

OpenStudy (anonymous):

no, it will be just three IBP

OpenStudy (anonymous):

Yep, what Jimmy said works, just do it reverse from what you usually do. Thanks Jimmy!

OpenStudy (anonymous):

but if u=x^(-3), then every time you differentiate u, the exponent gets farther from 0. So when do the IBPs stop?

OpenStudy (anonymous):

good point, i didnt think about that. :-)

OpenStudy (anonymous):

I think you are talaking about the tabular method in IBP , try using the direct theorem

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

What's the direct theorem?

OpenStudy (anonymous):

is it: uv-integral of(v(du))???

OpenStudy (anonymous):

But then again, the damn du/dx is 2, not 2x obviously. So you get pellet from IBP. And yes, you just add up on using u=x^-3... Direct theorem?

OpenStudy (anonymous):

yes. it is. @kumarsajan

OpenStudy (anonymous):

but when you use that it will make the du infinite if you chose x^-3 as your u, and if it is e^2x it is also infinite.

OpenStudy (anonymous):

I see !

OpenStudy (anonymous):

But, by using u=e^2x and dv=x^-3, you actually reduce x^-3 by integrating. So it eventually solves itself.

OpenStudy (anonymous):

thats the way: - u = e^2x

OpenStudy (anonymous):

i see then, as x becomes 0 it will be solved. :-)

OpenStudy (anonymous):

that's right